### v1

@@ 4.4 Global convergence
Suppose that $p_{\bm{x}}(\bm{\theta})$ is a member of the exponential family (11) with
$\bm{\eta}(\bm{\theta})=\Big{(}\sum\limits_{i\in[m]}r_{1,i}\ln(-g_{i}(\bm{\theta})),\ldots,\sum\limits_{i\in[m]}r_{p,i}\ln(-g_{i}(\bm{\theta}))\Big{)}^{\top}\in\mathbb{R}^{p}$ (47)
for some constants $r_{j,i}\in\mathbb{R},i\in[m];j\in[p]$. Suppose that a sequence of solutions, $\{\bm{\theta}^{(t)}\}_{t\geq 0}$, generated by (16) has the limit $\bm{\theta}^{*}$, and that for each $i\in[m]$,
$\lim_{t\to\infty}\frac{\sum_{j\in[p]}r_{j,i}\big{(}\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big{)}_{j}}{g_{i}(\bm{\theta}^{(t)})}\geq 0.$ (48)
Then $\bm{\theta}^{*}$ satisfies the Karush-Kuhn-Tucker condition for (46).
@@
5 Extended applications
For a matrix $\bm{H}\in\mathbb{R}^{p\times n}$, a vector $\bm{w}\in\mathbb{R}^{p}$, and a function $\phi$ on $\mathbb{R}^{p}$, consider
@@ 5 Extended applications
$\begin{array}[]{|lll}\underset{\bm{\theta}}{\mbox{minimize}}&f(\bm{\theta})=\phi(\bm{H}\bm{\theta}+\bm{w})\\ \mbox{subject to}&\bm{H}\bm{\theta}+\bm{w}\in\Theta,\end{array}~{}\mbox{or equivalently,}~{}~{}\begin{array}[]{|lll}\underset{\bm{\theta},\bm{\lambda}}{\mbox{minimize}}&\phi(\bm{\lambda})~{}\big{(}=f(\bm{\theta})\big{)}\\ \mbox{subject to}&\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},\bm{\lambda}\in\Theta.\end{array}$ (53)
@@
For $\bm{H}=\bm{I}$ and $\bm{w}=\bm{0}$, (53) is reduced to (1), and the examples presented in Section 4.2 can be equivalently cast as an example of this extension. For the extension, the condition (4) is modified. For (53), consider a distribution $\hat{p}_{\bm{x}}(\bm{\theta}):=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})$ satisfying $\mathrm{int}\Theta=\{\bm{\theta}\in\mathbb{R}^{n}\,|\,\hat{p}_{\bm{x}}(\bm{\theta})>0\}$. For such $\hat{p}_{\bm{x}}$, assume that there is a positive-valued function $G:\Xi\to(0,\infty)$ satisfying
$f(\bm{\theta})=-\ln\mathbb{E}_{\hat{p}_{\bm{x}}(\bm{\theta})}\big{[}G(\bm{X})\big{]}\quad\mbox{for all }\bm{\theta}\in\mathbb{R}^{n}\mbox{ such that }\hat{p}_{\bm{x}}(\bm{\theta})>0.$ (54)
@@
If we find such a tuple $(f,\Theta,\hat{p}_{\bm{x}},G)$, the EM scheme can be extended because $\hat{p}_{\bm{x}}(\bm{\theta})=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})=\beta(\bm{x})\exp\big{(}\bm{\eta}(\bm{H}\bm{\theta}+\bm{w})^{\top}\bm{\gamma}(\bm{x})-A(\bm{H}\bm{\theta}+\bm{w})\big{)}$ is a member of the exponential family, and the Fisher information matrix of $\hat{p}_{\bm{x}}(\bm{\theta})$ (or $p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})$) is given by $\bm{I}_{\hat{p}_{\bm{x}}}(\bm{\theta})=\bm{H}^{\top}\big{(}\mathbb{V}_{p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big{)}^{-1}\bm{H}.$
@@
In general, it is not possible to derive a simple natural gradient update formula for (53) like (25), (35), and (43), since the number of parameters is now less than that of random variables, and Theorem 3.3 cannot apply. However, the EM scheme is still applicable by minimizing a nonlinear convex optimization (16) at each iteration as far as an exponential family distribution with $A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})$ being convex is employed. Although the subproblem (16) might be as hard as (or harder than) the original problem, we may apply a generalized EM algorithm (GEM [10]), in which the M-step is terminated soon after a better incumbent is obtained. A straightforward example of such kind is the EM gradient algorithm [21], where a Newton method is applied to the subproblem. See Appendix C for the detailed descriptions.
5.1 Minimization of a polynomial over a polytope
For $\bm{B}\in\mathbb{R}^{(p-n)\times p}$, $\bm{c}\in\mathbb{R}^{p-n}$, $\bm{l},\bm{u}\in\mathbb{R}^{p}$ such that $\bm{u}-\bm{l}\in\mathbb{R}^{p}_{++}$, consider the generalized form of the problems (29) and (41):
@@ 5.1 Minimization of a polynomial over a polytope
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p}}{\mbox{minimize}}&F(\bm{\lambda})=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}\lambda_{j}^{n_{i,j}}\\ \mbox{subject to}&\bm{B}\bm{\lambda}=\bm{c},~{}\bm{l}\leq\bm{\lambda}\leq\bm{u}.\end{array}$ (57)
@@
Denote $\bm{B}=\left(\bm{B}_{1},\bm{B}_{2}\right)$ for $\bm{B}_{1}\in\mathbb{R}^{(p-n)\times n},\bm{B}_{2}\in\mathbb{R}^{(p-n)\times(p-n)}$, and assume that $\bm{B}_{2}$ is nonsingular. The dual form of Support Vector Classification can be of the form (57). The equality constraints, $\bm{B}\bm{\lambda}=\bm{c}$, of (57) can then be rewritten as $\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},$ where
@@
and (57) can be rewritten as
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p},\bm{\theta}\in\mathbb{R}^{n}}{\mbox{minimize}}&\phi(\bm{\lambda})=-\ln\Big{(}K-F\big{(}\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}+\bm{l}\big{)}\Big{)}\\ \mbox{subject to}&\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},~{}\bm{0}\leq\bm{\lambda}\leq\bm{1},\end{array}$ (60)
where $K$ is an upper bound of $F$ over the feasible set. Note that (32) is a special case of (60) with $(\bm{H},\bm{w},\bm{l},\bm{u})=(\bm{I},\bm{0},\bm{0},\bm{1})$. To apply the EM scheme to (60), let $\bm{X}=(X_{1},...,X_{p})$ be a vector of $p$ independent binomial random variables, $X_{j}\sim\mathrm{Bin}(m_{j},\lambda_{j})$, with $m_{j}:=\max\limits_{i\in[I]}n_{i,j},~{}j\in[p]$, so that
$p_{\bm{x}}(\bm{\lambda})$ $=\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}\lambda_{j}^{x_{j}}(1-\lambda_{j})^{m_{j}-x_{j}}=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})\equiv\hat{p}_{\bm{x}}(\bm{\theta}).$ (61)
Similar to Section 4.2, the condition (4) can be established.
@@
If $G$ is defined by
$G(\bm{x})$ $=K-\Big{(}\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}g_{i,j}(x_{j})+\tilde{b}\Big{)},$ (34)
where $g_{i,j}(x_{j})$ is defined as in Proposition 4.4 and $\tilde{a}_{i},~{}i\in[I]$, and $\tilde{b}$ are such that
$F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}+\bm{l})$ $\equiv\sum_{i\in[I]}a_{i}\prod_{j\in[p]}\big{(}(u_{j}-l_{j})\lambda_{j}+l_{j}\big{)}^{n_{i,j}}=\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}\lambda_{j}^{n_{i,j}}+\tilde{b},$
the condition (54) for (60) is fulfilled, namely, $f(\bm{\theta})(\equiv\phi(\bm{H}\bm{\theta}+\bm{w}))=-\ln\mathbb{E}_{\mathbb{P}}\big{[}G(\bm{X})\big{|}\bm{\theta}\big{]}$ holds with $\hat{p}_{\bm{x}}(\bm{\theta})$ defined by (61) and $\Theta=[\bm{0},\bm{1}]$, or $\phi(\bm{\lambda})=-\ln\mathbb{E}_{\mathbb{P}}\big{[}G(\bm{X})\big{|}\bm{\lambda}\big{]}$ holds with $p_{\bm{x}}(\bm{\lambda})$ defined by (33) and $\Theta=[\bm{0},\bm{1}]$.
From (14), the surrogate function $u(\cdot|\cdot)$ on $\mathrm{int}\Theta\times\mathrm{int}\Theta$ is a convex function of the form:
@@
where $\bm{h}_{j}$ denotes $j$-th row of $\bm{H}$, $\mathbb{E}_{\hat{q}_{\bm{x}}(\bar{\bm{\theta}})}[X_{j}]=\mathrm{diag}(\bar{\bm{\lambda}})\big{(}\bm{m}-\mathrm{diag}(\bm{1}-\bar{\bm{\lambda}})\,\nabla\phi(\bar{\bm{\lambda}})\big{)}.$ with $\bar{\bm{\lambda}}=\bm{H}\bar{\bm{\theta}}+\bm{w}$ via (15), and the EM algorithm for (60) is thus described as Algorithm 5. In general, however, simple update formula is no longer available unlike the examples in Section 4.
Algorithm 5 EM algorithm for (60)
1:Let $\bm{\theta}^{(0)}\in\mathbb{R}^{n}$ satisfy $\bm{l}<\bm{H}\bm{\theta}^{(0)}+\bm{w}<\bm{u}$, and set $t\leftarrow 1$.
@@
4:     $\mathbb{E}_{\hat{q}_{\bm{x}}(\bm{\theta}^{(t-1)})}[X_{j}]\leftarrow\mathrm{diag}(\bm{\lambda}^{(t-1)})\Big{(}\bm{m}-\mathrm{diag}\big{(}\bm{1}-\bm{\lambda}^{(t-1)}\big{)}\nabla\phi(\bm{\lambda}^{(t-1)})\Big{)}$,
@@
and the Sherman-Morrison-Woodbury formula yields the inverse of the Hessian as
@@
This can be beneficial for efficient implementation of the EM gradient when $p-n$ is small.
5.2 Quadratic program and linear program
Consider a quadratic program (QP):
$\begin{array}[]{|ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}&\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta}\\ \mbox{subject to}&\bm{A}^{\top}\bm{\theta}\leq\bm{c},\\ \end{array}$ (64)
where $\bm{A}\in\mathbb{R}^{n\times p},\bm{b}\in\mathbb{R}^{n},\bm{c}\in\mathbb{R}^{p}$, and $\bm{Q}\in\mathbb{S}^{p}$, so the objective can be nonconvex. It is easy to see that (64) falls in the form (53) with $\bm{\lambda}=\bm{c}-\bm{A}^{\top}\bm{\theta}$. Moreover, (64) includes a linear program (LP) as a special case by setting $\bm{Q}=\bm{O}$. We assume that (64) has an interior point.
To apply the EM algorithm to (64), consider $2p$ random variables $\bm{X}:=(\bm{X}_{1},\bm{X}_{2})$ such that $\bm{X}_{1}=(X_{1,1},...,X_{1,p})$ is $p$ independent Poisson random variables , $\bm{X}_{2}=(X_{2,1},...,X_{2,p})$ follows $p$-dimensional normal distribution $\bm{\mathrm{N}}(\bm{\theta}_{2},\bm{\Sigma})$ for $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$, and every pair of $X_{1,i}$ and $X_{2,j}$ is independent for $i,j\in[p]$. Specifically, its likelihood-like function of $\bm{X}=\bm{x}$ is given by
$p_{\bm{x}}(\bm{\theta})$ $=p_{\bm{x}_{1}}(\bm{\lambda})\,p_{\bm{x}_{2}}(\bm{\theta})=p_{\bm{x}_{1}}(\bm{c}-\bm{A}^{\top}\bm{\theta})\,p_{\bm{x}_{2}}(\bm{\theta})=\hat{p}_{\bm{x}_{1}}(\bm{\theta})\,p_{\bm{x}_{2}}(\bm{\theta}),$ (65)

### v2

@@ 4.4. Global convergence
Suppose that $p_{\bm{x}}(\bm{\theta})$ is a member of the exponential family (3.1) with
$\bm{\eta}(\bm{\theta})=\Big(\sum\limits_{i\in[M]}r_{1,i}\ln(-g_{i}(\bm{\theta})),\ldots,\sum\limits_{i\in[M]}r_{p,i}\ln(-g_{i}(\bm{\theta}))\Big)^{\top}\in\mathbb{R}^{p}$ (4.34)
for some constants $r_{j,i}\in\mathbb{R},i\in[M];j\in[p]$. Suppose that a sequence of solutions, $\{\bm{\theta}^{(t)}\}_{t\geq 0}$, generated by (3.6) has the limit $\bm{\theta}^{*}$, and that for each $i\in[M]$,
$\lim_{t\to\infty}\frac{\sum_{j\in[p]}r_{j,i}\big(\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t)})}[\bm{\gamma}(\bm{X})]-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]\big)_{j}}{g_{i}(\bm{\theta}^{(t)})}\geq 0.$ (4.35)
Then $\bm{\theta}^{*}$ satisfies the Karush–Kuhn–Tucker condition for (4.33).
@@
5. Extended applications
In this section, we consider extending the EM scheme for a more general optimization formulation of those examined in the previous section. Specifically, for a matrix $\bm{H}\in\mathbb{R}^{p\times n}$, a vector $\bm{w}\in\mathbb{R}^{p}$, and a function $\phi$ on $\mathbb{R}^{p}$, we consider
@@ 5. Extended applications
$\begin{array}[]{|lll}\underset{\bm{\theta}}{\mbox{minimize}}&f(\bm{\theta})=\phi(\bm{H}\bm{\theta}+\bm{w})\\ \mbox{subject to}&\bm{H}\bm{\theta}+\bm{w}\in\Theta,\end{array}~\mbox{or,}~~\begin{array}[]{|lll}\underset{\bm{\theta},\bm{\lambda}}{\mbox{minimize}}&\phi(\bm{\lambda})~\big(=f(\bm{\theta})\big)\\ \mbox{subject to}&\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},\bm{\lambda}\in\Theta.\end{array}$ (5.5)
@@
Two formulations in (5.5) are equivalent; for $\bm{H}=\bm{I}$ and $\bm{w}=\bm{0}$, they are both reduced to (2.1), and each of (4.2), (4.16), and (4.28) can be cast as a special case of this extension. However, to extend the EM scheme to problems of this general type (5.5), condition (2.4) needs to be modified. For (5.5), consider a distribution $\hat{p}_{\bm{x}}(\bm{\theta}):=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})$ that satisfies $\mathrm{int}\{\bm{\theta}\in\mathbb{R}^{n}|\bm{H}\bm{\theta}+\bm{w}\in\Theta\}=\{\bm{\theta}\in\mathbb{R}^{n}\,|\,\hat{p}_{\bm{x}}(\bm{\theta})>0\}$. For such $\hat{p}_{\bm{x}}$, assume that there is a positive-valued function $G:\Xi\to(0,\infty)$ satisfying
$f(\bm{\theta})=-\ln\mathbb{E}_{\hat{p}_{\bm{x}}(\bm{\theta})}\big[G(\bm{X})\big]\quad\mbox{for all }\bm{\theta}\in\mathbb{R}^{n}\mbox{ such that }\hat{p}_{\bm{x}}(\bm{\theta})>0.$ (5.6)
@@
If we find such a tuple $(f,\Theta,\hat{p}_{\bm{x}},G)$, the EM scheme can be extended because $\hat{p}_{\bm{x}}(\bm{\theta})=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})=\beta(\bm{x})\exp\big(\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{H}\bm{\theta}+\bm{w})-A(\bm{H}\bm{\theta}+\bm{w})\big)$ is a member of the exponential family, and the Fisher information matrix of $\hat{p}_{\bm{x}}(\bm{\theta})$ (or $p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})$) is given by $\bm{I}_{\hat{p}_{\bm{x}}}(\bm{\theta})=\bm{H}^{\top}\big(\mathbb{V}_{p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big)^{-1}\bm{H}.$
@@
In general, it is not possible to derive a simple natural gradient update formula for (5.5) like (4.6), (4.19), and (4.30), since the number of parameters is now less than that of random variables, and Theorem 3.3 cannot apply. However, the EM scheme is still applicable by minimizing a nonlinear convex optimization (3.6) at each iteration as far as an exponential family distribution with $A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})$ being convex is employed. Although the subproblem (3.6) might be as hard as (or harder than) the original problem, we may apply a generalized EM algorithm (GEM [13]), in which the M-step is terminated soon after a better incumbent is obtained. A straightforward example of such kind is the EM gradient algorithm [24], where a Newton method is applied to the subproblem. See Appendix C for the detailed descriptions.
5.1. Minimization of a polynomial over a polytope
The initial example of the extension (5.5) is the generalized form of the problems (4.10) and (4.28):
@@ 5.1. Minimization of a polynomial over a polytope
$\begin{array}[]{|ll}\underset{\bm{\nu}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&F(\bm{\nu}):=\sum\limits_{i\in[I]}a_{i}\prod\limits_{j\in[p]}\nu_{j}^{n_{i,j}}\\ \mbox{subject to}&\bm{B}\bm{\nu}=\bm{c},~\bm{l}\leq\bm{\nu}\leq\bm{u},\end{array}$ (5.9)
where $\bm{B}\in\mathbb{R}^{(p-n)\times p}$, $\bm{c}\in\mathbb{R}^{p-n}$ for $n,p\in\mathbb{N}$ such that $p>n$, and $\bm{l},\bm{u}\in\mathbb{R}^{p}$ such that $\bm{u}-\bm{l}\in\mathbb{R}^{p}_{++}$. We assume that (5.9) has an interior feasible point. For example, the dual form of Support Vector Classification can be of the form (5.9). We first show that if $\bm{B}$ has full row rank, i.e., $\mathrm{rank}\bm{B}=p-n$, (5.9) can be equivalently rewritten in the form of (5.5) with $\Theta=[\bm{0},\bm{1}]$ so that the EM scheme based on independent binomial distributions can be developed. Similarly to Section 4.2, the change of variables $\lambda_{j}=(\nu_{j}-l_{j})/(u_{j}-l_{j})$ for $j\in[p]$ yields an equivalent formulation to (5.9) as
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&\phi(\bm{\lambda})=-\ln\Big(K-F\big(\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}+\bm{l}\big)\Big)\\ \mbox{subject to}&\bm{B}\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}=\bm{c}-\bm{B}\bm{l},~\bm{0}\leq\bm{\lambda}\leq\bm{1}\end{array}$ (5.12)
@@
for an upper bound $K$ of $F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}+\bm{l})$ over the feasible region of (5.12). Without loss of generality, denoting $\bm{B}=\left(\bm{B}_{1},\bm{B}_{2}\right)$ for $\bm{B}_{1}\in\mathbb{R}^{(p-n)\times n},\bm{B}_{2}\in\mathbb{R}^{(p-n)\times(p-n)}$, we assume that $\bm{B}_{2}$ is nonsingular. The equality constraints of (5.12) can then be rewritten as $\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},$ where
@@
and (5.9) can be rewritten in the form of (5.5) with $\Theta=[\bm{0},\bm{1}]$, as
$\begin{array}[]{|ll}\underset{\bm{\lambda}\in\mathbb{R}^{p},\bm{\theta}\in\mathbb{R}^{n}}{\mbox{minimize}}\quad&\phi(\bm{\lambda})=-\ln\Big(K-\big(\sum\limits_{i\in[I]}\tilde{a}_{i}\prod\limits_{j\in[p]}\lambda_{j}^{n_{i,j}}+\tilde{b}\big)\Big)\\ \mbox{subject to}&\bm{\lambda}=\bm{H}\bm{\theta}+\bm{w},~\bm{0}\leq\bm{\lambda}\leq\bm{1},\end{array}$ (5.15)
for some $\tilde{a}_{i},~i\in[I]$, and $\tilde{b}$ such that
$F(\mathrm{diag}(\bm{u}-\bm{l})\bm{\lambda}+\bm{l})$ $=\sum_{i\in[I]}a_{i}\prod_{j\in[p]}\big((u_{j}-l_{j})\lambda_{j}+l_{j}\big)^{n_{i,j}}=\sum_{i\in[I]}\tilde{a}_{i}\prod_{j\in[p]}\lambda_{j}^{n_{i,j}}+\tilde{b}$.
To apply the EM scheme to (5.15), let $\bm{X}=(X_{1},...,X_{p})$ be a vector of $p$ independent binomial random variables, $X_{j}\sim\mathrm{Bin}(m_{j},\lambda_{j})$, with $m_{j}\geq\max\limits_{i\in[I]}n_{i,j},~j\in[p]$, so that
$p_{\bm{x}}(\bm{\lambda})$ $=\prod_{j\in[p]}\frac{m_{j}!}{x_{j}!(m_{j}-x_{j})!}\lambda_{j}^{x_{j}}(1-\lambda_{j})^{m_{j}-x_{j}}~\big(=\hat{p}_{\bm{x}}(\bm{\theta}):=p_{\bm{x}}(\bm{H}\bm{\theta}+\bm{w})\big).$ (5.16)
Similarly to Proposition 4.4 in Section 4.2, the following proposition shows a function $G(\cdot)$ that satisfies the modified condition (5.6) for problem (5.15) under the binomial probability distribution (5.16).
@@
Let $g_{i,j}(x_{j})$ be defined as in Proposition 4.4. If $G$ is defined by
[…]
the condition (5.6) for (5.15) is fulfilled, i.e., $f(\bm{\theta})(=\phi(\bm{H}\bm{\theta}+\bm{w}))=-\ln\mathbb{E}_{\mathbb{P}}\big[G(\bm{X})\big|\bm{\theta}\big]$ holds with $\hat{p}_{\bm{x}}(\bm{\theta})$ defined by (5.16) and $\Theta=[\bm{0},\bm{1}]$, or $\phi(\bm{\lambda})=-\ln\mathbb{E}_{\mathbb{P}}\big[G(\bm{X})\big|\bm{\lambda}\big]$ holds with $p_{\bm{x}}(\bm{\lambda})$ defined by (5.16) and $\Theta=[\bm{0},\bm{1}]$.
For $\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\{\bm{\theta}\in\mathbb{R}^{n}|\bm{H}\bm{\theta}+\bm{w}\in\Theta\}$, the surrogate function $u(\bm{\theta}|\bar{\bm{\theta}})$ is a convex function of the form:
@@
where $\bm{h}_{j}$ denotes $j$-th row of $\bm{H}$, $\mathbb{E}_{\hat{q}_{\bm{x}}(\bar{\bm{\theta}})}[\bm{X}]=\mathrm{diag}(\bar{\bm{\lambda}})\big(\bm{m}-\mathrm{diag}(\bm{1}-\bar{\bm{\lambda}})\,\nabla\phi(\bar{\bm{\lambda}})\big).$ with $\bar{\bm{\lambda}}=\bm{H}\bar{\bm{\theta}}+\bm{w}$ via (3.5), and the EM algorithm for (5.15) is thus described as Algorithm 5. In general, however, simple update formula is no longer available unlike the examples in Section 4.
Algorithm 5 EM algorithm for (5.15)
1:Let $\bm{\theta}^{(0)}\in\mathbb{R}^{n}$ satisfy $\bm{H}\bm{\theta}^{(0)}+\bm{w}\in(\bm{0},\bm{1})$, and set $t\leftarrow 1$.
@@
4:  $\mathbb{E}_{\hat{q}_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{X}]\leftarrow\mathrm{diag}(\bm{\lambda}^{(t-1)})\Big(\bm{m}-\mathrm{diag}\big(\bm{1}-\bm{\lambda}^{(t-1)}\big)\nabla\phi(\bm{\lambda}^{(t-1)})\Big)$,
@@
and the Sherman–Morrison–Woodbury formula yields the inverse of the Hessian as
@@
This can be beneficial for efficient implementation of the EM gradient when the size of $\bm{B}_{2}$, i.e., pnp-n is small.
5.2. Quadratic program and linear program
The second example of the extension (5.5) is a quadratic program (QP):
[…]
where $\bm{A}\in\mathbb{R}^{n\times p},\bm{b}\in\mathbb{R}^{n},\bm{c}\in\mathbb{R}^{p}$, and $\bm{Q}\in\mathbb{S}^{n}$, so the objective can be nonconvex. It is easy to see that (5.19) falls in the form (5.5) with $\bm{\lambda}=\bm{c}-\bm{A}^{\top}\bm{\theta}$ and $\Theta=\mathbb{R}^{p}_{+}$. Moreover, (5.19) includes a linear program (LP) as a special case by setting $\bm{Q}=\bm{O}$. We assume that (5.19) has an interior point.
To apply the EM algorithm to (5.19), consider $p+n$ random variables $\bm{X}:=(\bm{X}_{1},\bm{X}_{2})$ such that $\bm{X}_{1}=(X_{1,1},...,X_{1,p})$ is $p$ independent Poisson random variables, $\bm{X}_{2}=(X_{2,1},...,X_{2,n})$ follows $n$-dimensional normal distribution $\bm{\mathrm{N}}(\bm{\theta},\bm{\Sigma})$ for $\bm{\Sigma}\in\mathbb{S}^{n}_{++}$, and every pair of $X_{1,i}$ and $X_{2,j}$ is independent for $i\in[p],j\in[n]$. Specifically, its likelihood-like function of $\bm{X}=\bm{x}$ is given by
$p_{\bm{x}}(\bm{\theta})$ $=p_{\bm{x}_{1}}(\bm{\lambda})\,p_{\bm{x}_{2}}(\bm{\theta})=p_{\bm{x}_{1}}(\bm{c}-\bm{A}^{\top}\bm{\theta})\,p_{\bm{x}_{2}}(\bm{\theta})=\hat{p}_{\bm{x}_{1}}(\bm{\theta})\,p_{\bm{x}_{2}}(\bm{\theta}),$ (5.20)